Find three consecutive integers such that twice the smallest increased by the largest is 128.

thank you

Answer by stanbon(75887)

Find three consecutive integers such that twice the smallest increased by the largest is 128.
----------------------
1st: x-1
2nd: x
3rd: x+1
----------------
Equation:
2(x-1)+ (x+1) = 128
2x-2 + x + 1 = 128
3x = 129
x = 43
---
1st = x-1 = 42
2nd = x = 43
3rd = x+1 = 44
---------------------
Cheers,
Stan H.